OpenStudy (deadshot):

How do I solve 10^x = 1,000,000?

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf 10^x = 1,000,000 }\) \(\large\color{blue}{ \bf 10^x = 10^5 }\)

OpenStudy (anonymous):

first you should simplify it, to make it eaiser. try starting off with dividing both sides by a number both are divisible by.

OpenStudy (deadshot):

so then it would become \[1^x = 100,000\] when I divide both sides by 10, right?

OpenStudy (anonymous):

yes. now divide both sides by 1.

OpenStudy (mathstudent55):

Do what Solomon told you, but \(1,000,000 = 10^6\), not \(10^5\). Then use this: If \(10^x = 10^y\) then \(x = y\) If two powers are equal, and their bases are equal, then their exponents must be equal.

OpenStudy (mathstudent55):

Don't divide anything. Just do what I wrote above.

OpenStudy (anonymous):

so now the answer is, x=100,000 or 104 to the fourth power

OpenStudy (mathstudent55):

First step: write the given equation. Second step: rewrite the equation but change the right side to a power of 10. Third step: you have two powers of 10 equaling each other, write the exponents equal to each other. Fourth step. Done.

OpenStudy (anonymous):

if you divide, it just makes it simpler. #justsayin

OpenStudy (mathstudent55):

@JayeSoulei123 Once again, there is nothing to divide here by.

OpenStudy (deadshot):

@mathstudent55 so then if \[10^x=10^y\] , \[10^x=1,000,000\] , \[1,000,000 = 10^6\], and \[10^y = 10^6\] then \[10^x=10^6\]right?

OpenStudy (anonymous):

@mathstudent55 I didn't ask you for your opinion sir. go sit down yo.

OpenStudy (anonymous):

count the 0's that's x

OpenStudy (mathstudent55):

Ok, I'll spell it out. Here is the equation: 10^x = 1,000,000 Second line, change right side to power of 10. 10^x = 10^6 Third line: x = 6 Done

OpenStudy (deadshot):

Oh! Thanks! @mathstudent55

OpenStudy (anonymous):

why is he talking down now? he should for real swerve yo

OpenStudy (mathstudent55):

@JayeSoulei123 1. Anyone can give an opinion here. That's how OS works. 2. It's more helpful if you know what you're doing. When students need help, they need to learn how to solve a problem, not to be told incorrect information that will only confuse them more. 3. Perhaps I am just missing what you are trying to say, and it's my mistake. Then please write a complete solution to this problem and get to a final answer. If you know a method I don't know, I'd certainly like to learn it.

OpenStudy (anonymous):

oh shutup . @mathstudent55